# A Survey of Deep Graph Learning under Distribution Shifts: from Graph Out-of-Distribution Generalization to Adaptation

### v1

@@ 3.1 Model-Centric Approaches
Invariant Representation Learning. Invariant representation learning methods for graph OOD generalization aim to extract invariant information that is robust to distribution shifts across different environments [38, 39, 40]. Here, ”environments” refer to distinct settings that can influence the structure and relationships within the graph data. These methods can be further classified into (1) invariant subgraph extraction, (2) adversarial learning, and (3) disentangled representation learning.
@@
  • (1) Invariant Subgraph Extraction. Inspired by the invariant learning strategy in traditional OOD generalization tasks [41, 42, 43, 44], invariant subgraph extraction methods divide each graph into two segments [45, 46]: an invariant subgraph, which maintains stable structure-label relationships across distributions, and a variant subgraph, which may change across different distributions. By concentrating on the invariant subgraph, these methods leverage consistent relationships between graph structures and labels, which enhances the model’s ability to manage significant distribution shifts in OOD data, thereby improving its generalization performance. DIR [47] exemplifies this approach by distinguishing invariant subgraphs using a GNN-based approach, enabling the identification of consistent graph-label relations without the need for explicit distribution-based separation in the training data [47]. Conversely, GIL [45] advances this concept by utilizing the distributional divisions of training data to cluster variant subgraphs, thereby facilitating the generation of invariant graph representations for accurate classification. Additionally, MARIO [38] employs the graph Information Bottleneck (IB) principle [48, 49] for learning invariant representations by maximizing the mutual information between representations and labels while minimizing the mutual information between representations and inputs. Similarly, GSAT [46] formulates the graph IB by injecting learnable randomness into a node-level attention mechanism, which allows selecting subgraphs that keep invariant to the environmental shift and indicative to the labels.
@@
    Recent work by Piao et al. [50] propose a hierarchical approach for effective subgraph extraction. This method generates multiple semantic environments for each graph while ensuring consistency across hierarchical levels. By considering multiple environments and modeling the correlations between different environments within a hierarchical framework, this approach can better understand the inherent variability in the data, leading to improved robustness against distribution shifts. Sun et al. [51] propose DIVE framework, which trains multiple models on all label-predictive subgraphs, consisting of both spurious and invariant subgraphs. This encourages the models to focus on distinct structural patterns within these subgraphs, facilitating the precise extraction of invariant subgraphs and mitigating the effects of simplicity bias.
@@
  • (2) Adversarial Learning. Adversarial learning methods aim to learn invariant representations in different environments by enforcing models to perform consistently when adversarially enlarging the distribution shifts [52, 39]. Pioneering works of adversarial training adopt existing adversarial methods for OOD generalization from image data. For example, GNN-DRO [39] employs distributionally robust optimization [53] to address OOD challenges, aiming to minimize the worst-case loss within a defined range around the observed data distribution. However, the reliance of GNN-DRO on distribution labels can limit its application in the real world [54]. To address this limitation, GraphAT [40] introduces graph perturbations to create augmented distributions. These perturbations are achieved by maximizing the prediction discrepancies between a given node and its neighboring nodes, thereby effectively altering the graph’s smoothness to generate more diverse distributions.
  • (3) Disentangled Representation Learning. Disentangled representation learning aims to separate the distinct and informative factors in representations. In this manner, the learned disentangled representations contain decisive information for the target task. As a classic example, DisenGCN [55] introduces a multichannel convolutional layer along with a neighborhood routing strategy to achieve node representation disentanglement, and the disentangled representations are more robust to distribution shifts regarding node labels. IPGDN [56] enhances this disentanglement by promoting independence among latent factors based on Hilbert-Schmidt Independence Criterion (HSIC) [57], further improving generalization. From the perspective of factorization, FactorGCN [58] proposes to produce disentangled representations on the graph level by breaking down graphs into multiple interpretable factor graphs. As such, the graph-level generalization is strengthened. In addition to the graph-level generalization, NED-VAE [59] captures independent latent factors from both the node level and edge level with a variational autoencoder [60, 61]. This design improves generalization from various granularities of graph structures. Additionally, some works [62, 63, 64] point out that the degradation in OOD generalization ability often stems from spurious correlations between irrelevant patterns and category labels within the training data. To mitigate this, L2R-GNN [62] and OOD-GNN [63] utilize nonlinear graph decorrelation methods to eliminate spurious correlations between invariant patterns and other variables. L2R-GNN focuses on removing spurious correlations through a nonlinear approach, while OOD-GNN implements a global weight estimator for this purpose. Furthermore, EQuAD [65] proposes a graph invariance learning paradigm that disentangles invariant patterns from spurious features using the infomax principle [66, 67], thus enhancing the interpretability and robustness of learned representations.
    Beyond OOD generalization in conventional static graphs, dynamic graphs also face challenges from distribution shifts. Effectively modeling the complex latent environments and uncovering invariant patterns in dynamic graphs with distribution shifts remains an under-explored area. Zhang et al. propose DIDA [70], a framework that utilizes a disentangled spatio-temporal attention network to differentiate between variant and invariant patterns, thereby improving predictive stability in dynamic graphs under distribution shifts. I-DIDA [71] further refines predictive stability by identifying invariant patterns and applying disentangled spatio-temporal interventions. Similarly, Yuan et al. develop the EAGLE framework [72], which incorporates an environment-aware EA-DGNN to learn disentangled representations across various spatio-temporal environments. Zhou et al. [73] identify two key challenges in OOD generalization on spatiotemporal data: the entanglement of segment-level heterogeneity in temporal environments, and the failure of Invariant Risk Minimization (IRM) [42] on heterogeneous spatiotemporal observations. Their proposed CauSTG framework extracts invariant patterns through a spatiotemporal consistency learner and a hierarchical invariance explorer, directly addressing these challenges. Additionally, Wang et al. propose a STONE framework that employs a semantic graph learning module and a graph intervention mechanism capture invariant spatio-temporal dependencies among and thus address the spatio-temporal shifts [74].
Graph Self-supervised Learning. Graph self-supervised learning has gained significant attention for its role in graph OOD generalization by facilitating the learning of more generalizable representations. Specifically, graph self-supervised learning enhances generalizability by mapping similar graphs to close representations [78]. In this manner, graph self-supervised learning improves the GNN models’ understanding of new test graphs, as their representations can closely align with the training data [79]. Pretraining-GNN [80] is among the first to explore various graph self-supervised learning methods at both the node and graph levels to pre-train GNN models, including context prediction, attribute masking, and structural similarity prediction. To address graph size shifts, PATTERN [27] introduces a self-supervised task focused on learning d-pattern representations from local structures, thereby enhancing generalization from smaller to larger graphs. Moreover, OOD-GCL [81] resorts to graph contrastive learning on factorized graph representations, thereby excluding the information from the latent factors sensitive to distribution shifts and improving overall generalization.
More recently, various Graph Foundation Models (GFMs) [82] have also been proposed based on graph self-supervised learning. GFMs exhibit superior generalizability on a variety of downstream tasks. One of the pioneering works in this area, GPPT [83] generalizes GNN models with prompt-tuning. Specifically, based on pre-trained GNN models via self-supervised learning, GPPT introduces additional tokens that represent candidate labels, thereby improving generalization to new tasks without the tedious fine-tuning of GNNs. Based on the prompt-tuning strategy, GPF [84] proposes a universal prompt-tuning framework to pre-train GNN models across various self-supervised learning strategies. Operating on the input graph’s feature space, GPF improves the generalizability of the prompted graph to various downstream tasks in an adaptive way. Drawing inspiration from ControlNet [85], GraphControl [86] enhances generalization with a universal structural pre-training strategy that enables the alignment of the input space across distributions and incorporates characteristics of target distributions as conditional contexts.
@@
TABLE I: A summary of graph OOD generalization methods. ‘Task Level’ denotes the main task level, ‘Domain Source’ denotes whether the source domain consists of single or multiple domains, and ‘Distribution Shift’ denotes the distribution shifts that the methods aim to handle.
@@
3.2 Data-Centric Approaches
@@
Graph Data Augmentation. Graph data augmentation directly modifies graph structures and features to increase the diversity of existing training data, thereby exposing the model to a wider range of distributions to enhance its generalizability. Based on the used augmentation strategy, graph augmentation methods can be further classified into (1) graph prturbation and (2) graph mix-up.
  • (1) Graph Perturbation. For graph perturbation methods, original graphs are augmented by modifying the original graph structure such as adding or removing nodes and edges, as well as altering the representations of nodes and edges. In this manner, the augmented graphs are slightly different from the original graphs, thereby enhancing the diversity and quality of the training data. As a result, the model trained on the augmented data becomes more robust to various graph data distributions. Specifically, the goal of graph data augmentation methods for OOD generalization could be formulated as follows:
@@
    where $\textbf{A}^{\prime}$, $\textbf{X}^{\prime}$, and $\textbf{Y}^{\prime}$ represent the adjacency matrix, node feature matrix, and label of a perturbed graph, respectively. In this formulation, $g$ transforms the perturbed graph inputs into latent representations, while $f$ functions as the classifier.
@@
    As a classic example, Pattern-PT [27] studies the problem of graph size generalization and discovers that the discrepancy in small local structures across graphs could harm the generalizability of GNNs. To mitigate this issue, Pattern-PT proposes an interpolation strategy among graphs in order to learn comprehensive representations of local structures. P-gMPNN [93] focuses on the link-level OOD generalization problem with graph size shifts. This work explores the vulnerability of GNNs as link predictors based on structural node embeddings, and proposes a theoretically sound method that learns robust pairwise embeddings by modifying graphs into graphon fractions of common neighbors. To avoid causing significant changes and losing relevant information during augmentation, several works aim to enhance the stability of augmentation strategies. For instance, AIA [91] considers the causal invariance principle while enlarging the diversity of environments, aiming to reduce the distribution gap between the augmented environments and the potential test data unseen during training. To ensure causal invariance, AIA designs a causal generator and adversarially learns the masks that encapsulate causal information. Similarly, G-Splice [35] is proposed to avoid the corruption of the underlying causal information related to labels. The authors explore the extrapolation strategies from both the perspectives of structures and features to augment graph data without impacting the causal information. From the perspective of label shifts, LiSA [37] identifies that augmented graph labels may change during structural modifications, and hence proposes to avoid such a label shift issue during graph augmentation. Particularly, LiSA employs a variational subgraph generator, trained to produce label-invariant subgraphs, which are subsequently used in aiding graph OOD generalization. DLG [92] improves the label and distribution consistency by maximizing the mutual information between augmented and original graphs.
@@
  • (2) Graph Mix-up. The mixup strategies, initially proposed and shown to be effective in computer vision and natural language processing, have demonstrated significant advantages in enhancing generalization through interpolating instances [106, 107, 108]. These techniques work by creating synthetic training graph data through mix-up, which enhances model robustness to distribution shifts and shows promise in graph OOD generalization [109]. For instance, GraphMix [94] integrates manifold mixup [110] for generalizable node classification by simultaneously training a fully-connected network (FCN) and a GNN, with a parameter-sharing strategy that facilitates the transfer of essential node representations from the FCN to the GNN. Similarly, G-Mixup [95] extends this concept by interpolating latent node representations in the embedding space, offering generalizability across different graphs. Recognizing that latent representations may lack richness, NodeAug [111] and ifMixup [112] further propose to directly interpolate raw node features and graph data, respectively, to create more diverse inputs for GNNs.
@@
    On the other hand, $\mathcal{G}$-Mixup [96] proposes to perform mixup on graphons across classes in the Euclidean space, resulting in new graphs derived from the interpolation of graphons. The results demonstrate enhanced performance in graph classification tasks with distribution shifts and highlight the efficacy $\mathcal{G}$-mixup in promoting OOD generalization. Instead of relying on predefined disturbances, OOD-GMixup [97] generates virtual samples by applying manifold mixup directly to the graph rationale representations. This approach allows for the simulation of complex graph distribution shifts by manipulating these representations within a metric space. Furthermore, to accurately measure and control the training process amid these shifts, OOD-GMixup introduces an OOD confidence score, derived using Extreme Value Theory (EVT). This score, coupled with a sample re-weighting mechanism, actively manages the training dynamics by prioritizing virtual OOD samples, thereby enhancing the model’s generalization capabilities across varied and unforeseen graph distributions.
    While extensive research efforts [96, 94, 95] have focused on interpolation for synthetic data augmentation, relatively few studies have explored the extrapolation in addressing out-of-distribution challenges on graphs. Recently, a novel framework [113] was proposed to generate synthetic graphs using a non-Euclidean-space linear extrapolation method. The generated graphs have been theoretically validated to adhere to common causal assumptions. The framework consists of two components, which are designed to generate synthetic graphs in both the structural and feature spaces. Extensive experiments demonstrate that this extrapolation approach enhances a model’s OOD generalization ability on graphs and outperforms previous invariant learning methods [45, 47, 11] and other graph data augmentation techniques [96, 94, 95].
@@
Distribution Augmentation. In contrast to graph data augmentation, which enhances the original graph at the graph level, distribution augmentation focuses on generating new distributions of graphs to facilitate robust representation learning [54]. Distribution augmentation is particularly useful when the distribution available during training is scarce, making learning robust representations extremely challenging, as the model struggles to distinguish between decisive information and spurious information that varies across distributions.
@@
GREA [98] is a pioneering augmentation approach to compensate for the lack of sufficient distribution. It operates under the assumption of the graph rationale, i.e., for each input graph, there exists a specific subgraph that is decisive for its prediction. Therefore, GREA proposes to augment each graph by replacing the variant subgraph (i.e., subgraphs that do not involve the graph rationale) of the target graph with the variant subgraphs from other existing graphs. In this way, by training on these augmented graphs with varying domain subgraphs, the models will learn to focus solely on the graph rationale for making inferences, thereby enhancing OOD generalization. On the other hand, DPS [100] develops and utilizes several subgraph generators to generate diverse augmented subgraphs as additional distribution, while preserving the subgraphs with graph rationale. An additional regularization term is further introduced to amplify the difference among augmented distributions, enhancing the diversity of distributions of model training in order to improve the generalizability of the model. Following this, IGM [103] identifies invariant subgraphs and generates diverse environments by concatenating these with environment-related subgraphs. This targeted mixing approach not only helps eliminate spurious correlations but also fosters sufficient distribution shifts essential for robust invariant learning.
@@
Additionally, EERM [54] proposes to generate new distribution by leveraging invariance principles for prediction. EERM employs multiple context explorers, namely the graph structure editors, which are adversarially trained to maximize the variance of risks across multiple generated distributions. Such a design helps address the challenge of generating new distribution caused by the inter-connections among nodes on graphs. Based on EERM, FLOOD [99] further incorporates a bootstrapped learning module that could be refined on the test data to improve OOD generalization. ERASE [102] enhances generalization by employing a decoupled label propagation approach that corrects noisy labels through structural denoising before training. This method ensures that the model learns from more accurate representations of the data, effectively distinguishing meaningful information from noise. By maximizing coding rate reduction during training, ERASE adapts to diverse data distributions, improving robustness against label noise.
4 Training-Time Graph OOD Adaptation

### v2

@@ III-A Model-Centric Approaches
Invariant Representation Learning. Invariant representation learning methods for graph OOD generalization aim to extract invariant information that is robust to distribution shifts across different environments [38, 39, 40]. Here, ”environments” refer to distinct settings that can influence the structure and relationships within the graph data. These methods can be further classified into (1) invariant subgraph extraction, (2) adversarial learning, and (3) disentangled representation learning. These subcategories reflect specific learning techniques to learn invariant features from graph data. To provide a clearer understanding of these methods, we illustrate their key ideas in Figure 3.
@@
  • (1) Invariant Subgraph Extraction. Inspired by the invariant learning strategy in traditional OOD generalization tasks [41, 42, 43, 44], invariant subgraph extraction methods divide each graph into two distinct segments [45, 46]: an invariant subgraph, which maintains stable structure-label relationships across different distributions, and a variant subgraph, which may change across distributions. By concentrating on the invariant subgraph, these methods enable models to learn representations that levearge stable graph-label dependencies rather than distribution-specific variations, thereby improving its generalization performance. For example, DIR [47] exemplifies this approach by distinguishing invariant subgraphs using a GNN-based approach, enabling the identification of consistent graph-label relations without the need for explicit distribution-based separation in the training data [47]. Conversely, GIL [45] advances this concept by utilizing the distributional divisions of training data to cluster variant subgraphs, thereby facilitating the generation of invariant graph representations for accurate classification. Additionally, GSAT [46] formulates the graph Information Bottleneck (IB) principle [48, 49] by injecting learnable randomness into a node-level attention mechanism, which allows selecting subgraphs that keep invariant to the environmental shift and indicative to the labels.
@@
    A recent work by Piao et al. [50] proposes a hierarchical approach for effective subgraph extraction. This method constructs multiple semantic environments for each graph while ensuring consistency across hierarchical levels. By considering multiple environments and modeling the correlations between different environments within a hierarchical framework, this approach can better understand the inherent variability in the data, leading to improved robustness against distribution shifts. Sun et al. [51] propose DIVE framework, which trains multiple models on all label-predictive subgraphs, consisting of both spurious and invariant subgraphs. This encourages the models to focus on distinct structural patterns within these subgraphs, facilitating the precise extraction of invariant subgraphs and mitigating the effects of simplicity bias.
    Figure 3: Illustration of three invariant representation learning approaches: (1) Invariant Subgraph Extraction partitions graph structures for stability, (2) Adversarial Learning enhances robustness via perturbations, and (3) Disentangled Representation Learning separates independent factors for better generalization.
@@
  • (2) Adversarial Learning. Adversarial learning methods aim to learn invariant representations in different environments by enforcing models to perform consistently under adversarially induced distribution shifts [52, 39]. These methods leverage adversarial perturbations or optimization strategies to strengthen the model’s ability to handle unseen distribution changes. Pioneering works in adversarial training have adapted existing adversarial techniques from image-based OOD generalization to graph learning. For example, GNN-DRO [39] employs distributionally robust optimization [53] to address OOD challenges, aiming to minimize the worst-case loss within a defined range around the observed data distribution. However, the reliance of GNN-DRO on distribution labels can limit its application in the real world [54]. To address this limitation, GraphAT [40] introduces graph perturbations to create augmented distributions. These perturbations are achieved by maximizing the prediction discrepancies between a given node and its neighboring nodes, thereby effectively altering the graph’s smoothness to generate more diverse distributions.
@@
    TABLE II: A summary of graph OOD generalization methods. ‘Task Level’ denotes the main task level, ‘Domain Source’ denotes whether the source domain consists of single or multiple domains, and ‘Distribution Shift’ denotes the distribution shifts that the methods aim to handle.
@@
  • (3) Disentangled Representation Learning. Disentangled representation learning aims to separate the distinct factors of variation within the data, where each factor captures a specific aspect of the graph structure. By isolating these independent factors, the approach ensures that they remain stable across different distribution shifts. This disentanglement allows the model to focus on the most invariant and task-relevant features of the data, thereby improving generalization across various domains or environments. As a classic example, DisenGCN [58] introduces a multichannel convolutional layer along with a neighborhood routing strategy to achieve node representation disentanglement, and the disentangled representations are more robust to distribution shifts regarding node labels. IPGDN [59] enhances this disentanglement by promoting independence among latent factors based on Hilbert-Schmidt Independence Criterion (HSIC) [93], further improving generalization. From the perspective of factorization, FactorGCN [60] proposes to produce disentangled representations on the graph level by breaking down graphs into multiple interpretable factor graphs. As such, the graph-level generalization is strengthened. In addition to the graph-level generalization, NED-VAE [61] captures independent latent factors from both the node level and edge level with a variational autoencoder [94, 95]. This design improves generalization from various granularities of graph structures. Additionally, some works [65, 66, 74] point out that the degradation in OOD generalization ability often stems from spurious correlations between irrelevant patterns and category labels within the training data. To mitigate this, L2R-GNN [65] and OOD-GNN [66] utilize nonlinear graph decorrelation methods to eliminate spurious correlations between invariant patterns and other variables. L2R-GNN focuses on removing spurious correlations through a nonlinear approach, while OOD-GNN implements a global weight estimator for this purpose. Furthermore, EQuAD [67] proposes a graph invariance learning paradigm that disentangles invariant patterns from spurious features using the infomax principle [96, 97], thus enhancing the interpretability and robustness of learned representations.
    Beyond OOD generalization in conventional static graphs, dynamic graphs also face challenges from distribution shifts. Effectively modeling the complex latent environments and uncovering invariant patterns in dynamic graphs with distribution shifts remains an under-explored area. Zhang et al. propose DIDA [68], a framework that utilizes a disentangled spatio-temporal attention network to differentiate between variant and invariant patterns, thereby improving predictive stability in dynamic graphs under distribution shifts. I-DIDA [64] further refines predictive stability by identifying invariant patterns and applying disentangled spatio-temporal interventions. Similarly, Yuan et al. develop the EAGLE framework [69], which incorporates an environment-aware EA-DGNN to learn disentangled representations across various spatio-temporal environments. Zhou et al. [70] identify two key challenges in OOD generalization on spatiotemporal data: the entanglement of segment-level heterogeneity in temporal environments, and the failure of Invariant Risk Minimization (IRM) [42] on heterogeneous spatiotemporal observations. Their proposed CauSTG framework extracts invariant patterns through a spatiotemporal consistency learner and a hierarchical invariance explorer, directly addressing these challenges.
Graph Self-supervised Learning. Graph self-supervised learning (SSL) has emerged as an effective strategy for improving graph OOD generalization by enabling models to learn transferable representations without relying on labeled data. The key objective is to enforce consistency in learned representations across different views or perturbations of the graph, ensuring that similar graphs are mapped to close representations [99]. This enhances the model’s ability to generalize to unseen test graphs by capturing structural and semantic patterns that remain stable across distribution shifts [7]. Pretraining-GNN [75] is among the first to explore various graph self-supervised learning methods at both the node and graph levels to pre-train GNN models, including context prediction, attribute masking, and structural similarity prediction. To address graph size shifts, PATTERN [27] introduces a self-supervised task focused on learning d-pattern representations from local structures, thereby enhancing generalization from smaller to larger graphs. Moreover, OOD-GCL [76] resorts to graph contrastive learning on factorized graph representations, thereby excluding the information from the latent factors sensitive to distribution shifts and improving overall generalization.
More recently, various Graph Foundation Models (GFMs)[100] have been developed based on graph self-supervised learning, demonstrating strong generalizability across diverse downstream tasks. One of the pioneering works in this area, GPPT[77], extends GNN models with prompt-tuning. Specifically, GPPT builds upon pre-trained GNNs trained via self-supervised learning and introduces additional tokens representing candidate labels, enabling better generalization to new tasks without requiring extensive fine-tuning. Based on the prompt-tuning strategy, GPF [78] proposes a universal prompt-tuning framework to pre-train GNN models across various self-supervised learning strategies. By operating in the feature space of input graphs, GPF improves the generalizability of the prompted graph to various downstream tasks in an adaptive way. Drawing inspiration from ControlNet [101], GraphControl [79] enhances generalization with a universal structural pre-training strategy that enables the alignment of the input space across distributions and incorporates characteristics of target distributions as conditional contexts.
III-B Data-Centric Approaches
@@
Graph Data Augmentation. Graph data augmentation directly modifies the structures and features of the graphs within the training set. This increases the diversity of the data, exposing the model to various graph topologies, node features, and edge relationships. By simulating variations of the input graphs, augmentation techniques allow the model to learn more robust and generalizable representations. Based on the augmentation strategy, graph augmentation methods can be further classified into (1) graph perturbation and (2) graph mix-up.
  • (1) Graph Perturbation. Graph perturbation methods involve small, controlled modifications to graph structures, such as adding or removing edges, or slightly altering node features, which introduces variability into the training data while preserving the original graph’s core characteristics. The augmented graphs, although slightly different from the originals, provide new and diverse training examples that capture the variability that the model might encounter in real-world applications. By increasing the diversity of the graphs in the training set, graph perturbation methods help the model become more robust to a range of graph data distributions. Specifically, the goal of graph data augmentation methods for OOD generalization could be formulated as follows:
@@
    where $\textbf{A}^{\prime}$, $\textbf{X}^{\prime}$, and $\textbf{Y}^{\prime}$ represent the adjacency matrix, node feature matrix, and label of a perturbed graph, respectively. In this formulation, $g$ transforms the perturbed graph inputs into latent representations, while $f$ serves as the classifier.
@@
    As a classic example, Pattern-PT [27] studies the problem of graph size generalization and discovers discrepancies in small local structures across graphs can negatively impact the generalizability of GNNs. To mitigate this issue, Pattern-PT proposes an interpolation strategy among graphs in order to learn comprehensive representations of local structures. P-gMPNN [82] focuses on the link-level OOD generalization problem under graph size shifts. This work explores the vulnerability of GNNs as link predictors based on structural node embeddings, and proposes a theoretically sound method that learns robust pairwise embeddings by modifying graphs into graphon fractions of common neighbors. To prevent excessive modifications that might lead to information loss, several works aim to enhance the stability of augmentation strategies. For instance, AIA [80] considers the causal invariance principle while enlarging the diversity of environments, aiming to reduce the distribution gap between the augmented environments and the potential test data unseen during training. To ensure causal invariance, AIA designs a causal generator and adversarially learns the masks that encapsulate causal information. Similarly, G-Splice [35] is proposed to avoid the corruption of the underlying causal information related to labels. The authors explore the extrapolation strategies from both the perspectives of structures and features to augment graph data without impacting the causal information. From a label-shift perspective, LiSA [37] identifies that augmented graph labels may change during structural modifications, and hence proposes to avoid such a label shift issue during graph augmentation. Particularly, LiSA employs a variational subgraph generator, trained to produce label-invariant subgraphs, which are subsequently used in aiding graph OOD generalization. DLG [81] improves the label and distribution consistency by maximizing the mutual information between augmented and original graphs.
@@
  • (2) Graph Mix-up. The mix-up strategies, initially proposed and shown to be effective in computer vision and natural language processing, have demonstrated significant advantages in enhancing generalization through interpolating instances [104, 105, 106]. These techniques generate synthetic training graph data through mix-up, which enhances model robustness to distribution shifts and shows promise in graph OOD generalization [107]. For instance, GraphMix [83] integrates manifold mixup [108] for generalizable node classification by simultaneously training a fully-connected network (FCN) and a GNN, with a parameter-sharing strategy that facilitates the transfer of essential node representations from the FCN to the GNN. Similarly, G-Mixup [84] extends this approach by interpolating latent node representations in the embedding space, offering generalizability across different graphs. Since latent representations may not fully capture the richness of the original data, NodeAug [109] and ifMixup [110] further propose to directly interpolate raw node features and graph data, respectively, to create more diverse inputs for GNNs.
@@
    On the other hand, $\mathcal{G}$-Mixup [85] proposes to perform mixup on graphons across classes in the Euclidean space, resulting in new graphs derived from the interpolation of graphons. The results demonstrate enhanced performance in graph classification tasks with distribution shifts and highlight the efficacy in promoting OOD generalization. Instead of relying on predefined disturbances, OOD-GMixup [86] generates virtual samples by applying manifold mixup directly to the graph rationale representations. This approach enables the simulation of complex graph distribution shifts by manipulating these representations within a metric space. Furthermore, to accurately measure and control the training process amid these shifts, OOD-GMixup introduces an OOD confidence score, derived using Extreme Value Theory (EVT). This score, coupled with a sample re-weighting mechanism, actively manages the training dynamics by prioritizing virtual OOD samples, thereby enhancing the model’s generalization capabilities across varied and previously unseen graph distributions. While extensive research efforts [85, 83, 84] have focused on interpolation for synthetic data augmentation, relatively few studies have explored the extrapolation in addressing out-of-distribution challenges on graphs. Recently, a novel framework [111] was proposed to generate synthetic graphs using a non-Euclidean-space linear extrapolation method. The generated graphs have been theoretically validated to adhere to common causal assumptions. The framework consists of two components, which are designed to generate synthetic graphs in both the structural and feature spaces. Extensive experiments demonstrate that this extrapolation approach enhances a model’s OOD generalization ability on graphs and outperforms previous invariant learning methods [45, 47, 11] as well as existing graph data augmentation techniques [85, 83, 84].
@@
Distribution Augmentation. In contrast to graph data augmentation, which enhances the original graph at the graph level, distribution augmentation generates new distributions of graphs to facilitate robust representation learning [54]. This approach is particularly beneficial when the available training distribution is limited, making it difficult for the model to learn robust representations. In such cases, the model may struggle to distinguish between decisive information and spurious information that vary across distributions.
@@
One of the pioneering methods in this domain is GREA [87], which is designed to address the challenge of scarce and limited graph distributions. GREA operates under the assumption that each input graph contains a specific subgraph, known as the ”graph rationale,” that is crucial for making accurate predictions. To leverage this, GREA augments graphs by replacing variant subgraphs (i.e., those not part of the graph rationale) in a target graph with variant subgraphs from other graphs. In this way, by training on these augmented graphs with varying domain subgraphs, the model will learn to focus solely on the graph rationale for making inferences, thereby enhancing OOD generalization. Similarly, DPS [89] generates diverse augmented subgraphs while preserving those containing the graph rationale. Multiple subgraph generators are used to produce additional distributions that capture different variations of the graph. To further improve the diversity of these augmented distributions, DPS incorporates a regularization term that increases the variance between them, improving generalizability and robustness in the face of distributional changes. Following this, IGM [92] identifies invariant subgraphs and generates diverse environments by concatenating these with environment-related subgraphs. This co-mixup strategy not only creates diverse environments by mixing variant subgraphs, but also refines learning by mixing invariant subgraphs to eliminate spurious correlations and enhance OOD generalization.
@@
Additionally, EERM [54] introduces a distribution augmentation approach based on invariance principles. This approach leverages multiple context explorers, specifically graph structure editors, which are adversarially trained to maximize the variance in risk across the generated graph distributions. This design is particularly effective in dealing with the inherent complexity of graph structures, where node interconnections complicate direct augmentation. Building on EERM, FLOOD [88] further incorporates a bootstrapped learning module that could be refined on the test data to improve OOD generalization. ERASE [91] enhances generalization by employing a decoupled label propagation approach that corrects noisy labels through structural denoising before training. This method ensures that the model learns from more accurate representations of the data, effectively distinguishing meaningful information from noise. By maximizing coding rate reduction during training, ERASE adapts to diverse data distributions, improving robustness against label noise.
IV Training-Time Graph OOD Adaptation